Some students planned a get together. The budget for the food was $500. Five of the students failed to come because of the distance and therefore the cost of food for each member increased by $5. How many students attended the get-together?
Answer by stanbon(57361)

Some students planned a get together. The budget for the food was $500. Five of the students failed to come because of the distance and therefore the cost of food for each member increased by $5. How many students attended the get-together?
-----------
Let number of students be "x".
Budget was $500 or average of $500/x per student
-----------
Five no-shows leaves "x-5" students with average of $500/(x-5)
-------------
EQUATION:
new average - old average = $5
500/(x-5) - (500/x) = 5
Multiply thru by x(x-5) to get:
500x - 500(x-5) = 5x(x-5)
500x - 500x+ 2500 = 5x^2-25x
5x^2-25x-2500 = 0
x^2-5x-500=0
x = [5+-sqrt(5^2-4*-500)]/2
Positive answer only:
x = [5+sqrt(2025)]/2
x = [5+45]/2
x = 25 students
=================
Cheers,
Stan H.